OpenStudy (anonymous):

2 questions guys. Need Answers Quick Please! 1) Write (cos theta -1)/(sin theta) in terms of a single trigonometric function a)tan theta/2 b)-tan theta/2 c)tan theta d)-tan theta 2)Write 1 - 2sin^2(2theta) in terms of a single trigonometric function a)sin4theta b)sin2theta c)cos4theta d)cos2theta Thanks!

OpenStudy (vishweshshrimali5):

I sincerely hope this is not some test/quiz problem. We cannot give you direct answers.

OpenStudy (vishweshshrimali5):

But that doesn't mean I can't give you hints :)

OpenStudy (vishweshshrimali5):

For problem 1: Use this: \[\large{cos(\theta) = 1 - 2\sin^2{\cfrac{\theta}{2}}}\]

OpenStudy (vishweshshrimali5):

For problem 2: Use the one I mentioned in hint for problem 1

OpenStudy (anonymous):

I think i got number 2. Is it Cos4theta? and im still lost with the first problem

OpenStudy (anonymous):

1st one is easier, with the hint given.

OpenStudy (anonymous):

Is the second one right? Can you walk me through the first one?

OpenStudy (anonymous):

it's identities. Do you know about it ?

OpenStudy (anonymous):

ya which identity? the chapter i studied only had double angle identities, power reducing identities and half angle identities

OpenStudy (anonymous):

I'll give you all the formulas required, you solve it, if you can't. I'll help you then. :)

OpenStudy (anonymous):

K thanks!

OpenStudy (anonymous):